Prove that the points A(-4,-1), B(-2,-3), C(4,3), and D(2,5) are the vertices of a rectangle. Sketch the rectangle.
Answer by Alan3354(69443)

Prove that the points A(-4,-1), B(-2,-3), C(4,3), and D(2,5) are the vertices of a rectangle. Sketch the rectangle.
---------------
Find the slope of AB, BC, CD and AD
If the lines are perpendicular, it's a rectangle.
----
For AB:
slope m = diffy/diffx = -2/2 = -1
--
For BC:
m = 6/6 = 1
--
The 2 slopes are negative reciprocals, so they're perpendicular.
-------------------------
Do the same for the others.